OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). (1 points) Question 9 options: 1) -9x + 27 2) -9x + 25 3) -9x + 5 4) -9x - 2

zepdrix (zepdrix):

\[\large \color{royalblue}{g(x)=-x+3}\]\[\large f\left[\color{orangered}{x}\right]=9\color{orangered}{x}-2\] So we have a composition of functions. We want to plug our function g(x) into f. \[\large f\left[\color{royalblue}{g(x)}\right]=9\color{royalblue}{g(x)}-2\]Plugging in our g(x) gives us,\[\large f\left[\color{royalblue}{g(x)}\right]=9\color{royalblue}{(-x+3)}-2\]Do you think you can simplify it from there? Lemme know if you have any questions.

OpenStudy (anonymous):

thanks its b i also i have 2 more questions Let f(x) = x - 3 and g(x) = x + 11. Find f(x) • g(x). (1 points) Question 4 options: 1) x2 - 8x - 33 2) x2 - 8x + 14 3) x2 + 8x + 14 4) x2 + 8x - 33

OpenStudy (anonymous):

Let f(x) = x2 + 3x - 4 and g(x) = x + 5. Find f(x) • g(x). (1 points) Question 5 options: 1) x3 + 3x2 + 16x - 20 2) x3 + 5x2 + 14x - 20 3) x3 + 8x2 + 11x - 20 4) x3 + 9x2 + 19x - 20